# A Randomized Method for Simulating Lindblad Equations and Thermal State Preparation

### v4

@@ I Introduction
The open quantum system is a fundamental research area in quantum physics, focusing on the behavior of quantum systems interacting with external environments [1]. In contrast to closed quantum systems, which are isolated and unaffected by their surroundings, open quantum systems experience complex dynamics due to external influences, resulting in phenomena such as decoherence and dissipation. Central to this topic is the Lindblad master equation [2, 3], which is the Markovian approximation of a general open quantum dynamics that models the system evolution through dissipative quantum jumps. The Lindblad dynamics has nowadays found fruitful applications in various fields, including materials science [4], quantum optics [5], and quantum computation [6]. In particular, recent developments employed Lindblad dynamics as an algorithmic tool for quantum Gibbs sampling [7, 8, 9, 10, 11], aiming at preparing the thermal or ground state on a quantum computer.
@@
Given the wide applications of Lindblad dynamics, developing efficient simulation methods for Lindblad equations is of great importance. Over the past decade, both classical and quantum algorithms have been extensively explored for this purpose. Classical methods [12] often face challenges due to exponential scaling in complexity as the system size increases. In contrast, quantum algorithms show promise in reducing this cost from exponential to polynomial scaling. However, many existing quantum methods [13, 14, 15] face challenges such as requiring a significant number of ancilla qubits and complex control circuits, especially when dealing with Lindblad equations that involve a large number of jump operators.
@@
In this paper, we propose a randomized quantum algorithm for simulating the Lindblad equation. Instead of encoding all jump operators into a quantum circuit, our method implements a single randomly sampled jump operator at each time step. This approach allows for a simpler Lindbladian simulation with the complexity independent of the number of jump operators. This simplification is particularly advantageous when dealing with a large or infinite number of jump operators. Our method draws inspiration from the qDRIFT approach for Hamiltonian simulation [16], which we generalize here from closed to open quantum systems.
@@
Our approach is particularly useful for quantum Gibbs sampling, which is analogous to classical Markov chain Monte Carlo methods. The Lindblad dynamics-based quantum Gibbs sampler aims to construct an ergodic quantum Markov dynamics that converges to the thermal state. Existing approaches [7, 9, 10] typically rely on deterministic methods, requiring a large number of jump operators to ensure ergodicity. In contrast, our method utilizes randomly sampled jump operators at each step, leveraging randomness to simplify the implementation.
@@
The rigorous convergence analysis will be provided in Section III. The randomized approximation (3) reduces the task of simulating an ensemble of Lindbladians to simulating a single Lindbladian $\mathcal{L}_{a}$ with simple $H_{a}$ and $V_{a}$ over short time intervals. This enables a simpler algorithmic implementation for each time step. For example, we can dilate the Lindbladian into a Hamiltonian acting on the system coupled with an ancilla qubit and then implement Hamiltonian simulation on the joint system [8].
I.1 Overview of Main Results
I.1.1 Analysis of randomized Lindbladian simulation
@@
These two setups are motivated by the literature on the qDRFIT method for Hamiltonian simulation [16, 17]. In terms of observables, the error bound for the average channel (5) controls the difference between the expectation values obtained from the randomized simulation algorithm and those of the actual dynamics, i.e., $|\operatorname{\mathbb{E}}\langle O\rangle_{\mathcal{E}_{\tau,M}(\rho)}-\langle O\rangle_{e^{T\bar{\mathcal{L}}}(\rho)}|$, where $\langle O\rangle_{\rho}:={\mathrm{Tr}}(O\rho)$ denotes the expectation of an observable $O$ with respect to a state $\rho$. The analysis of typical realizations of the random channel (6) further provides bounds on the deviation from the expected behavior $\operatorname{\mathbb{E}}|\langle O\rangle_{\mathcal{E}_{\tau,M}(\rho)}-\langle O\rangle_{e^{T\bar{\mathcal{L}}}(\rho)}|$; see Corollaries 5 and 8.
Assuming that the approximation error of the numerical scheme $\mathcal{F}_{\tau}(\mathcal{L}_{a})$ to $e^{\tau\mathcal{L}_{a}}$ is $\mathcal{O}(\tau^{2})$ (see remarks after Assumption 2), we provide a non-asymptotic error analysis of the randomized algorithm (4) for simulating Lindblad dynamics in both the average and random setting. We first consider the convergence in the average sense. Let $T=M\tau$ be the total evolution time. Theorem 4 shows that the error $\|e^{T\bar{\mathcal{L}}}-\bar{\mathcal{E}}_{\tau,M}\|_{\diamond}$ between the ideal evolution and the averaged implementation (5), in terms of the diamond distance (7), scales as $\mathcal{O}(T\tau)=\mathcal{O}(T^{2}/M)$, matching the error bound of the qDRIFT protocol given in [16] for Hamiltonian simulation.
Then, we study the convergence of the typical realization (6) with respect to a weighted $\ell^{2}$-metric (10), which is a variant of the standard quantum $\chi^{2}$-divergence [18]. We show in Theorem 7 that the averaged weighted $\ell^{2}$-error of simulating (1) via the random channel (6) for time $T=M\tau$ generally scales as $\mathcal{O}(\sqrt{T\tau})=\mathcal{O}(T/\sqrt{M})$. This result generalizes the analysis of the random product formulas for Hamiltonian systems [19, 17] to open systems. It is noteworthy that we employ the weighted $\ell^{2}$-norm here rather than the more commonly used trace norm in the error analysis of quantum simulations. The reason for this choice will be discussed in Remark 2.
I.1.2 Applications to quantum Gibbs sampling
@@
A notable example is the Clifford-random Davies generator, which employs an ensemble of jump operators with uncorrelated random Hermitian coupling matrices (see 6). This construction enables efficient algorithmic implementation while guaranteeing rapid mixing for certain classes of models.
From a computational perspective, the random coupling matrix can be implemented using a unitary design, which admits efficient realization through random Clifford circuits [20]. This randomization approach enables the simulation of Lindbladian dynamics with exponentially many jump operators. It is also worth noting that, compared to the implementation of a general Davies generator [7], this random construction allows us to bypass the need for isolating Bohr frequencies by simulating the Heisenberg evolution of coupling operators for an infinitely long period and simultaneous block-encoding techniques.
Regarding convergence properties, the Clifford-random Davies generator induces a classical random walk on the spectral domain, leading to a mixing time bound determined exclusively by the spectral density of the system. For strongly interacting Hamiltonians obeying the semicircle law, the spectral gap of this random Davies generator scales as $\Omega(\beta^{-3/2})$, where $\beta$ denotes the inverse temperature. This behavior is particularly relevant for systems described by random matrix theory, including SYK models of interacting fermions [21, 22] and nuclear energy level statistics [23], where such spectral characteristics are observed.
Our result provides the first instance of a fast-converging quantum Gibbs sampling algorithm in the strongly interacting regime, extending previous works that primarily focused on weakly interacting systems [24, 25, 26]. This advancement broadens the literature on efficient thermalization in quantum Gibbs sampling.
We note that [27] also provides a fast thermalization analysis for ergodic open quantum systems by reducing the dynamics to a classical random walk on the spectrum. Their analysis is based on the assumption that the coupling operator satisfies the Eigenstate Thermalization Hypothesis (ETH). In contrast, we construct a coupling operator using random Hermitian matrices generated through Clifford random gates, which results in a simpler analysis and enables a concrete algorithmic implementation of the dynamics.
II Notations and Preliminaries
@@
The Lindblad dynamics (also called quantum Markov semigroup) is a $C_{0}$-semigroup of quantum channels $(e^{t\mathcal{L}})_{t\geq 0}$, which is governed by the Lindblad master equation with the following GKSL form [28, 29]:
@@
For self-adjoint $X\in\mathcal{B}(\mathcal{H})$, the variance of $X$ with respect to $\sigma\in\mathcal{D}_{+}(\mathcal{H})$ is defined as [30, Section 2.2]
@@
where $P_{I}^{\perp}$ denote the projection onto $\mathrm{span}\{{I}\}^{\perp}$ with respect to the KMS norm. This leads to the definition of quantum $\chi^{2}$-divergence between two quantum states $\rho$ and $\sigma$ [18]:
@@
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\chi^{2}(\rho,\sigma)=\|\rho-\sigma\|_{\sigma,-1/2}^{2}$}\,.
The metric $\|\cdot\|_{\sigma,-1/2}$ shall be used for measuring errors in the analysis of the random channel (6). Finally, we remark that both the quantum $\chi^{2}$-divergence and the weighted $\ell^{2}$-norm can be framed within the broader context of monotone Riemannian metrics in quantum information theory [31, 32]. Specifically, a Riemannian metric in this setting is a positive definite bilinear form $M_{\sigma}(A,B)$ with $\sigma\in\mathcal{D}_{+}(\mathcal{H})$, defined for traceless Hermitian tangent operators $A,B$ in the space $\mathcal{T}_{P}$, given by
@@
As shown in [31, 32], there exists a one-to-one correspondence between the family of monotone Riemannian metrics and a class of convex operator functions. Of particular relevance to our work is the following special one:
@@
It is clear that $\chi^{2}$-divergence is recovered as $\chi^{2}(\rho,\sigma)=M_{\sigma}(\rho-\sigma,\rho-\sigma)$, and, upon relaxing the traceless condition for $A,B$, $M_{\sigma}$ yields the weighted $\ell^{2}$-metric (10). The monotonicity of $M_{\sigma}$ readily implies the contractivity of the weighted norm $\|\cdot\|_{\sigma,-1/2}$ under any quantum channel $\Phi$ (cf. [18, Theorem 4]):
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\Phi(Y)\|_{\Phi(\sigma),-1/2}\leq\|Y\|_{\sigma,-1/2}.$} (11)
@@
We refer to [18, 33] for a detailed discussion on the detailed balance condition and the spectral gap.
III Convergence Analysis for Lindbladian Simulation
@@
III.1 Analysis of the Average Channel
@@
We refer to [14, Section 4] for the proof of Lemma 3. This bound is crucial for the analysis of the approximation error of the random product approximation (3). We also make the following assumption that the local truncation error of implementing a single Lindbladian is of order $\mathcal{O}(\tau^{2})$.
@@
Let $\{\mathcal{L}_{a}\}_{a\in\mathcal{A}}$ satisfy 1. For the quantum simulation $\mathcal{F}_{\tau}(\mathcal{L}_{a})$ of a Lindblad evolution $e^{\tau\mathcal{L}_{a}}$ up to a small time $\tau=\mathcal{O}(\lambda^{-1})$, we assume
@@
We will demonstrate in Appendix B that the condition (12) holds for the Hamiltonian-based algorithm presented by [14, 8]. It is noteworthy that higher-order methods $\mathcal{F}_{\tau}$ are not helpful here. Even if the single-time-step evolution $e^{\tau\mathcal{L}_{a}}$ is implemented exactly, the qDRIFT approximation (3) still results in an $\mathcal{O}(\tau^{2})$ local truncation error, since the product (3) is essentially a randomized compilation of the first-order Lie-Trotter decomposition. Consequently, a first-order quantum simulation for $e^{\tau\mathcal{L}_{a}}$, as proposed in [14, 8], is sufficient for our purposes. Additional examples of first-order methods for $e^{\tau\mathcal{L}_{a}}$ can be found in [12, 34].
@@
The proof is given in Appendix C. We outline the proof idea below. From 2, the local truncation error of the average channel $\bar{\mathcal{E}}_{\tau,M}$ is $\mathcal{O}(\lambda^{2}\tau^{2})$. Combining this with the contractivity of the quantum channel, we know that the error accumulates linearly over time. Thus, we obtain an $M\mathcal{O}(\lambda^{2}\tau^{2})=\mathcal{O}(\lambda^{2}T\tau)=\mathcal{O}(\lambda^{2}T^{2}/M)$ error bound. As an immediate corollary of Theorem 4, we obtain the following error bound for observable estimation.
@@
The Trotter-Suzuki product-formula algorithm is widely recognized as a simple yet efficient method for simulating Hamiltonian dynamics on a quantum computer; see [35] and references therein for recent developments. For simplicity, let $\mathcal{A}$ be a discrete ensemble with a uniform distribution $\mu$. Then, the dynamics in (1) reduce to the standard case with a finite number of jump operators:
@@
where we assume that $\mathcal{L}_{a}$ is sufficiently simple, in the sense that simulating $e^{\tau\mathcal{L}_{a}}$ over a short time $\tau$ has a gate complexity of $\mathcal{O}(1)$. Kliesch et al. [36] introduced the first quantum algorithm for simulating Lindblad dynamics, based on the first-order Lie-Trotter formula, with a gate complexity of $\mathcal{O}\left(|\mathcal{A}|\lambda^{2}T^{2}/\epsilon\right)$ for (13) under Assumption 1. In contrast, our randomized approach reduces the complexity to $\mathcal{O}\left(\lambda^{2}T^{2}/\epsilon\right)$, making it independent of the number of jump operators. Thus, when $|\mathcal{A}|$ is large, as is often the case in applications in thermal state preparation or electronic structure Hamiltonians, our method could substantially reduce the simulation cost.
The scaling in simulation time $T$ and accuracy $\epsilon$ can be further improved using the second-order Suzuki formula, which achieves a gate complexity of $\mathcal{O}(|\mathcal{A}|\sqrt{\lambda^{3}T^{3}/\epsilon})$, as shown in [37]. However, unlike Hamiltonian simulation, higher-order product formulas cannot be directly applied to Lindblad dynamics [13]. This limitation arises because, unlike unitary dynamics, Markovian quantum dynamics cannot evolve backward in time while obtaining a CPTP map. More recent works [13, 14, 15, 34] have significantly improved simulation costs using techniques such as Stinespring dilation, oblivious amplitude amplification, and Dyson series expansions. These methods could achieve costs that scale linearly with $T$ and polylogarithmically with $\epsilon$, but they still depend at least linearly on $|\mathcal{A}|$.
@@
After completing the first manuscript, we became aware of concurrent works aimed at improving lower-order product-formula-based Lindbladian simulation [38, 39, 40]. Specifically, Borras and Marvian [38] proposed a quantum algorithm that combines randomized compiling with the second-order product formula, achieving a gate complexity of $\mathcal{O}(\sqrt{\lambda^{3}T^{3}/\epsilon})$. However, their method is limited to a special class of Lindbladians where the short-time evolution of the dissipative part is assumed to be approximately a convex combination of unitary channels with a fixed state. The work by Peng et al. [39] introduced a new approximation channel for short-time Lindblad evolution, motivated by the classical quantum trajectory method. They designed two quantum algorithms with reduced dependence on the number of jump operators. Notably, their first algorithm [39, Theorem 1] achieves the same scaling as ours, but with an additional dependence on the system size $n$. Moreover, David et al. [40] extended the randomized approach from Hamiltonian simulation [41] to Lindblad dynamics, proposing first- and second-order randomized Trotter-Suzuki formulas that achieve gate complexities of $\mathcal{O}(|\mathcal{A}|\sqrt{\lambda^{3}T^{3}/\epsilon})$ and $\mathcal{O}(\sqrt{|\mathcal{A}|\lambda^{3}T^{3}/\epsilon})$, respectively. The key idea is to randomize the order of the summands in the product formula, which is qualitatively different from the qDRIFT method. Additionally, in [40, Section 5], a qDRIFT-type algorithm is proposed with the same error bound as ours (Theorem 4). However, in addition to this result, we also establish the error for each individual realization (see Section III.2). Beyond these algorithmic advances, another main contribution of our work lies in applying the proposed randomized framework to the quantum thermal state preparation (see Section IV).
III.2 Analysis of the Random Channel

### v5

@@ 1 Introduction
The open quantum system is a fundamental research area in quantum physics, focusing on the behavior of quantum systems interacting with external environments [breuer2007theory]. In contrast to closed quantum systems, which are isolated and unaffected by their surroundings, open quantum systems experience complex dynamics due to external influences, resulting in phenomena such as decoherence and dissipation. Central to this topic is the Lindblad master equation [cmp/1103899849, cmp/1103860160], which is the Markovian approximation of a general open quantum dynamics that models the system evolution through dissipative quantum jumps. The Lindblad dynamics has nowadays found fruitful applications in various fields, including materials science [Harbola_2006], quantum optics [doi:10.1080/00107514.2021.1924281], and quantum computation [PhysRevLett.75.3788]. In particular, recent developments employed Lindblad dynamics as an algorithmic tool for quantum Gibbs sampling [chen2023quantum, ding2023singleancilla, chen2023efficient, ding2024efficient, gilyén2024quantum], aiming at preparing the thermal or ground state on a quantum computer.
@@
Given the wide applications of Lindblad dynamics, developing efficient simulation methods for Lindblad equations is of great importance. Over the past decade, both classical and quantum algorithms have been extensively explored for this purpose. Classical methods [cao2021structure] often face challenges due to exponential scaling in complexity as the system size increases. In contrast, quantum algorithms show promise in reducing this cost from exponential to polynomial scaling. However, many existing quantum methods [childs2016efficient, cleve2016efficient, li2023simulating] face challenges such as requiring a significant number of ancilla qubits and complex control circuits, especially when dealing with Lindblad equations that involve a large number of jump operators.
@@
In this paper, we propose a randomized quantum algorithm for simulating the Lindblad equation. Instead of encoding all jump operators into a quantum circuit, our method implements a single randomly sampled jump operator at each time step. This approach allows for a simpler Lindbladian simulation with the complexity independent of the number of jump operators. This simplification is particularly advantageous when dealing with a large or infinite number of jump operators. Our method draws inspiration from the qDRIFT approach for Hamiltonian simulation [Campbell_2019], which we generalize here from closed to open quantum systems.
@@
Our approach is particularly useful for quantum Gibbs sampling, which is analogous to classical Markov chain Monte Carlo methods. The Lindblad dynamics-based quantum Gibbs sampler aims to construct an ergodic quantum Markov dynamics that converges to the thermal state. Existing approaches [chen2023quantum, chen2023efficient, ding2024efficient] typically rely on deterministic methods, requiring a large number of jump operators to ensure ergodicity. In contrast, our method utilizes randomly sampled jump operators at each step, leveraging randomness to simplify the implementation.
@@
The rigorous convergence analysis will be provided in Section 3. The randomized approximation (3) reduces the task of simulating an ensemble of Lindbladians to simulating a single Lindbladian $\mathcal{L}_{a}$ with simple $H_{a}$ and $V_{a}$ over short time intervals. This enables a simpler algorithmic implementation for each time step. For example, we can dilate the Lindbladian into a Hamiltonian acting on the system coupled with an ancilla qubit and then implement Hamiltonian simulation on the joint system [ding2023singleancilla].
1.1 Overview of Main Results
1.1.1 Analysis of randomized Lindbladian simulation
@@
These two setups are motivated by the literature on the qDRFIT method for Hamiltonian simulation [Campbell_2019, Chen_2021]. In terms of observables, the error bound for the average channel (5) controls the difference between the expectation values obtained from the randomized simulation algorithm and those of the actual dynamics, i.e., $|\operatorname{\mathbb{E}}\langle O\rangle_{\mathcal{E}_{\tau,M}(\rho)}-\langle O\rangle_{e^{T\bar{\mathcal{L}}}(\rho)}|$, where $\langle O\rangle_{\rho}:={\mathrm{Tr}}(O\rho)$ denotes the expectation of an observable $O$ with respect to a state $\rho$. The analysis of typical realizations of the random channel (6) further provides bounds on the deviation from the expected behavior $\operatorname{\mathbb{E}}|\langle O\rangle_{\mathcal{E}_{\tau,M}(\rho)}-\langle O\rangle_{e^{T\bar{\mathcal{L}}}(\rho)}|$; See Corollary 5 and Corollary 8.
Assuming that the approximation error of the numerical scheme $\mathcal{F}_{\tau}(\mathcal{L}_{a})$ to $e^{\tau\mathcal{L}_{a}}$ is $\mathcal{O}(\tau^{2})$ (see remarks after Assumption 2), we provide a non-asymptotic error analysis of the randomized algorithm (4) for simulating Lindblad dynamics in both the average and random setting. We first consider the convergence in the average sense. Let $T=M\tau$ be the total evolution time. Theorem 4 shows that the error $\|e^{T\bar{\mathcal{L}}}-\bar{\mathcal{E}}_{\tau,M}\|_{\diamond}$ between the ideal evolution and the averaged implementation (5), in terms of the diamond distance (7), scales as $\mathcal{O}(T\tau)=\mathcal{O}(T^{2}/M)$, matching the error bound of the qDRIFT protocol given in [Campbell_2019] for Hamiltonian simulation.
Then, we study the convergence of the typical realization (6) with respect to a weighted $\ell^{2}$-metric (10), which is a variant of the standard quantum $\chi^{2}$-divergence [temme2010chi]. We show in Theorem 7 that the averaged weighted $\ell^{2}$-error of simulating (1) via the random channel (6) for time $T=M\tau$ generally scales as $\mathcal{O}(\sqrt{T\tau})=\mathcal{O}(T/\sqrt{M})$. This result generalizes the analysis of the random product formulas for Hamiltonian systems [huang2020matrix, Chen_2021] to open systems. It is noteworthy that we employ the weighted $\ell^{2}$-norm here rather than the more commonly used trace norm in the error analysis of quantum simulations. The reason for this choice will be discussed in Remark 2.
1.1.2 Applications to quantum Gibbs sampling
@@
A notable example is the Clifford-random Davies generator, which employs an ensemble of jump operators with uncorrelated random Hermitian coupling matrices (see Assumption 6). This construction enables efficient algorithmic implementation while guaranteeing rapid mixing for certain classes of models.
From a computational perspective, the random coupling matrix can be implemented using a unitary design, which admits efficient realization through random Clifford circuits [Dankert2009]. This randomization approach enables the simulation of Lindbladian dynamics with exponentially many jump operators. It is also worth noting that, compared to the implementation of a general Davies generator [chen2023quantum], this random construction allows us to bypass the need for isolating Bohr frequencies by simulating the Heisenberg evolution of coupling operators for an infinitely long period and simultaneous block-encoding techniques.
Regarding convergence properties, the Clifford-random Davies generator induces a classical random walk on the spectral domain, leading to a mixing time bound determined exclusively by the spectral density of the system. For strongly interacting Hamiltonians obeying the semicircle law, the spectral gap of this random Davies generator scales as $\Omega(\beta^{-3/2})$, where $\beta$ denotes the inverse temperature. This behavior is particularly relevant for systems described by random matrix theory, including SYK models of interacting fermions [Kitaev2015, feng2019spectrum] and nuclear energy level statistics [PhysRevLett.52.1], where such spectral characteristics are observed.
Our result provides the first instance of a fast-converging quantum Gibbs sampling algorithm in the strongly interacting regime, extending previous works that primarily focused on weakly interacting systems [kochanowski2024rapid, rouze2024efficient, bergamaschi2024quantum]. This advancement broadens the literature on efficient thermalization in quantum Gibbs sampling.
We note that [chen2023fastthermalizationeigenstatethermalization] also provides a fast thermalization analysis for ergodic open quantum systems by reducing the dynamics to a classical random walk on the spectrum. Their analysis is based on the assumption that the coupling operator satisfies the Eigenstate Thermalization Hypothesis (ETH). In contrast, we construct a coupling operator using random Hermitian matrices generated through Clifford random gates, which results in a simpler analysis and enables a concrete algorithmic implementation of the dynamics.
2 Notations and Preliminaries
@@
The Lindblad dynamics (also called quantum Markov semigroup) is a $C_{0}$-semigroup of quantum channels $(e^{t\mathcal{L}})_{t\geq 0}$, which is governed by the Lindblad master equation with the following GKSL form [gorini1976completely, lindblad1976generators]:
@@
For self-adjoint $X\in\mathcal{B}(\mathcal{H})$, the variance of $X$ with respect to $\sigma\in\mathcal{D}_{+}(\mathcal{H})$ is defined as [muller2018sandwiched, Section 2.2]
@@
where $P_{I}^{\perp}$ denote the projection onto $\mathrm{span}\{{I}\}^{\perp}$ with respect to the KMS norm. This leads to the definition of quantum $\chi^{2}$-divergence between two quantum states $\rho$ and $\sigma$ [temme2010chi]:
@@
$\chi^{2}(\rho,\sigma)=\|\rho-\sigma\|_{\sigma,-1/2}^{2}$}\,.
The metric $\|\cdot\|_{\sigma,-1/2}$ shall be used for measuring errors in the analysis of the random channel (6). Finally, we remark that both the quantum $\chi^{2}$-divergence and the weighted $\ell^{2}$-norm can be framed within the broader context of monotone Riemannian metrics in quantum information theory [petz1996geometries, lesniewski1999monotone]. Specifically, a Riemannian metric in this setting is a positive definite bilinear form $M_{\sigma}(A,B)$ with $\sigma\in\mathcal{D}_{+}(\mathcal{H})$, defined for traceless Hermitian tangent operators $A,B$ in the space $\mathcal{T}_{P}$, given by
@@
As shown in [petz1996geometries, lesniewski1999monotone], there exists a one-to-one correspondence between the family of monotone Riemannian metrics and a class of convex operator functions. Of particular relevance to our work is the following special one:
@@
It is clear that $\chi^{2}$-divergence is recovered as $\chi^{2}(\rho,\sigma)=M_{\sigma}(\rho-\sigma,\rho-\sigma)$, and, upon relaxing the traceless condition for $A,B$, $M_{\sigma}$ yields the weighted $\ell^{2}$-metric (10). The monotonicity of $M_{\sigma}$ readily implies the contractivity of the weighted norm $\|\cdot\|_{\sigma,-1/2}$ under any quantum channel $\Phi$ (cf. [temme2010chi, Theorem 4]):
$\|\Phi(Y)\|_{\Phi(\sigma),-1/2}\leq\|Y\|_{\sigma,-1/2}.$} (11)
@@
We refer to [temme2010chi, Carlen_2019] for a detailed discussion on the detailed balance condition and the spectral gap.
3 Convergence Analysis for Lindbladian Simulation
@@
3.1 Analysis of the Average Channel
@@
We refer to [cleve2016efficient, Section 4] for the proof of Lemma 3. This bound is crucial for the analysis of the approximation error of the random product approximation (3). We also make the following assumption that the local truncation error of implementing a single Lindbladian is of order $\mathcal{O}(\tau^{2})$.
@@
Let $\{\mathcal{L}_{a}\}_{a\in\mathcal{A}}$ satisfy Assumption 1. For the quantum simulation $\mathcal{F}_{\tau}(\mathcal{L}_{a})$ of a Lindblad evolution $e^{\tau\mathcal{L}_{a}}$ up to a small time $\tau=\mathcal{O}(\lambda^{-1})$, we assume
@@
We will demonstrate in Appendix B that the condition (12) holds for the Hamiltonian-based algorithm presented by [cleve2016efficient, ding2023singleancilla]. It is noteworthy that higher-order methods $\mathcal{F}_{\tau}$ are not helpful here. Even if the single-time-step evolution $e^{\tau\mathcal{L}_{a}}$ is implemented exactly, the qDRIFT approximation (3) still results in an $\mathcal{O}(\tau^{2})$ local truncation error, since the product (3) is essentially a randomized compilation of the first-order Lie-Trotter decomposition. Consequently, a first-order quantum simulation for $e^{\tau\mathcal{L}_{a}}$, as proposed in [cleve2016efficient, ding2023singleancilla], is sufficient for our purposes. Additional examples of first-order methods for $e^{\tau\mathcal{L}_{a}}$ can be found in [cao2021structure, ding2024simulating].
@@
The proof is given in Appendix LABEL:sec:prof1. We outline the proof idea below. From Assumption 2, the local truncation error of the average channel $\bar{\mathcal{E}}_{\tau,M}$ is $\mathcal{O}(\lambda^{2}\tau^{2})$. Combining this with the contractivity of the quantum channel, we know that the error accumulates linearly over time. Thus, we obtain an $M\mathcal{O}(\lambda^{2}\tau^{2})=\mathcal{O}(\lambda^{2}T\tau)=\mathcal{O}(\lambda^{2}T^{2}/M)$ error bound. As an immediate corollary of Theorem 4, we obtain the following error bound for observable estimation.
@@
The Trotter-Suzuki product-formula algorithm is widely recognized as a simple yet efficient method for simulating Hamiltonian dynamics on a quantum computer; see [Childs_2021] and references therein for recent developments. For simplicity, let $\mathcal{A}$ be a discrete ensemble with a uniform distribution $\mu$. Then, the dynamics in (1) reduce to the standard case with a finite number of jump operators:
@@
where we assume that $\mathcal{L}_{a}$ is sufficiently simple, in the sense that simulating $e^{\tau\mathcal{L}_{a}}$ over a short time $\tau$ has a gate complexity of $\mathcal{O}(1)$. Kliesch et al. [Kliesch_2011] introduced the first quantum algorithm for simulating Lindblad dynamics, based on the first-order Lie-Trotter formula, with a gate complexity of $\mathcal{O}\left(|\mathcal{A}|\lambda^{2}T^{2}/\epsilon\right)$ for (13) under Assumption 1. In contrast, our randomized approach reduces the complexity to $\mathcal{O}\left(\lambda^{2}T^{2}/\epsilon\right)$, making it independent of the number of jump operators. Thus, when $|\mathcal{A}|$ is large, as is often the case in applications in thermal state preparation or electronic structure Hamiltonians, our method could substantially reduce the simulation cost.
The scaling in simulation time $T$ and accuracy $\epsilon$ can be further improved using the second-order Suzuki formula, which achieves a gate complexity of $\mathcal{O}(|\mathcal{A}|\sqrt{\lambda^{3}T^{3}/\epsilon})$, as shown in [werner2016positive]. However, unlike Hamiltonian simulation, higher-order product formulas cannot be directly applied to Lindblad dynamics [childs2016efficient]. This limitation arises because, unlike unitary dynamics, Markovian quantum dynamics cannot evolve backward in time while obtaining a CPTP map. More recent works [childs2016efficient, cleve2016efficient, li2023simulating, ding2024simulating] have significantly improved simulation costs using techniques such as Stinespring dilation, oblivious amplitude amplification, and Dyson series expansions. These methods could achieve costs that scale linearly with $T$ and polylogarithmically with $\epsilon$, but they still depend at least linearly on $|\mathcal{A}|$.
@@
After completing the first manuscript, we became aware of concurrent works aimed at improving lower-order product-formula-based Lindbladian simulation [borras2025quantum, peng2024quantum, david2024faster]. Specifically, Borras and Marvian [borras2025quantum] proposed a quantum algorithm that combines randomized compiling with the second-order product formula, achieving a gate complexity of $\mathcal{O}(\sqrt{\lambda^{3}T^{3}/\epsilon})$. However, their method is limited to a special class of Lindbladians where the short-time evolution of the dissipative part is assumed to be approximately a convex combination of unitary channels with a fixed state. The work by Peng et al. [peng2024quantum] introduced a new approximation channel for short-time Lindblad evolution, motivated by the classical quantum trajectory method. They designed two quantum algorithms with reduced dependence on the number of jump operators. Notably, their first algorithm [peng2024quantum, Theorem 1] achieves the same scaling as ours, but with an additional dependence on the system size $n$. Moreover, David et al. [david2024faster] extended the randomized approach from Hamiltonian simulation [childs2019faster] to Lindblad dynamics, proposing first- and second-order randomized Trotter-Suzuki formulas that achieve gate complexities of $\mathcal{O}(|\mathcal{A}|\sqrt{\lambda^{3}T^{3}/\epsilon})$ and $\mathcal{O}(\sqrt{|\mathcal{A}|\lambda^{3}T^{3}/\epsilon})$, respectively. The key idea is to randomize the order of the summands in the product formula, which is qualitatively different from the qDRIFT method. Additionally, in [david2024faster, Section 5], a qDRIFT-type algorithm is proposed with the same error bound as ours (Theorem 4). However, in addition to this result, we also establish the error for each individual realization (see Section 3.2). Beyond these algorithmic advances, another main contribution of our work lies in applying the proposed randomized framework to the quantum thermal state preparation (see Section 4).
3.2 Analysis of the Random Channel